OpenStudy (anonymous):

Pleaseee help ! The area of a square is 128 m^2 (to the 2nd power). How long is t diagonal?

Parth (parthkohli):

@satellite73 Congratulations for one year of OS, in advance!

OpenStudy (anonymous):

if the area is 128 then the length of the side is \(\sqrt{128}\) and so the length of the diagonal is \(\sqrt{128}\times \sqrt{2}=\sqrt{256}\)

OpenStudy (anonymous):

??

Parth (parthkohli):

Your profile says "Days Studied: 364"

Parth (parthkohli):

One year of logging in*

OpenStudy (anonymous):

@BamaBamaa it is also true that \(\sqrt{256}=16\)

OpenStudy (anonymous):

@ParthKohli no kidding? i will have to check